# Towards Designing a Question-Answering Chatbot for Online News: Understanding Perspectives and Questions

### 17 Dec 2023 v1

@@ 5.2.4. Linguistic Features
This figure shows multiple visual elements and mapping between them.
Table 3. Results of the SEM model. Each cell represents the coefficient for the respective relation. p-value significance: * p ¡ 0.05, ** p ¡ 0.01, *** p ¡ 0.001.
Authors Chatbot
Fact
-0.15*
0.06
-0.03
-0.03
-0.14
0.11
-0.08
-0.03
Details
0.50***
-0.07
-0.09
0.18
0.41**
-0.14
-0.02
0.17
Evidence
-0.49***
0.11
001
-0.11
0.10
-0.05
-0.02
0.16
Closed-ended
0.05
-0.16
-0.10
0.25
0.18*
0.02
0.03
-0.09
Opinion
0.23**
-0.009
0.09
-0.09
0.1
-0.09
0.05
0.06
Explanation
-0.28**
0.02
0.13
-0.15
0.1
-0.18
0.06
0.17
Journalistic Integrity
-0.61***
0.14
0.02
-0.16
0.10
-0.08
0.001
0.02
Bloom’s Taxonomy
0.06
0.27
-0.06
-0.20
0.14*
-0.07
0.43
-0.1
Length
15.3
22.21
3.76
30.62
11.12
16.66
46.75
6.45
Sentiment
-0.03
0.01
-0.02
0.02
-0.10
0.01
0.00
0.01
Table 3. Results of the SEM model. Each cell represents the coefficient for the respective relation. p-value significance: * p ¡ 0.05, ** p ¡ 0.01, *** p ¡ 0.001.
@@ 5.3. Effect of Perceived Quality and Personal Preferences (RQ3, RQ4)
To measure how perceived quality influences questions, we adopted a structural equation model (SEM) (Ullman and Bentler, 2012), shown in Figure 5. Similar to previous research (Graefe et al., 2018), we consider perceived quality to be a latent variable and derive it from user-provided ratings on readability, credibility, and journalistic expertise. Figure 9 shows participants’ ratings across the three higher-level dimensions that capture the perceived quality of the articles. Overall, participants rated the articles highly across the three dimensions. To answer RQ4, we also added participants’ preferences for news outlets in the model. We manually checked the preferred news outlets provided by the participants (Figure 7e), matched the polarity of the outlets to AllSides ratings (AllSides, 2023), and then categorized them to either left, neutral, or right-leaning news outlets. We assigned neutral to participants who read both left and right-leaning news (27). We assigned a left-leaning tag if the participant consumed mostly left-leaning news (60) and right-leaning (37) vice versa. Finally, we fit two separate datasets to the model, one for the author’s condition and one for the chatbot condition.
Table 3 presents the results from SEM. The Root Mean Square Error of Approximation (RMSEA) values for the two models are 0.06 and 0.07, indicating good model fits. The coefficients show the polarity and strength of the relations. For example, the coefficient for Journalistic Integrity and Perceived Quality was -0.41. This means low-quality articles correlated with a high number of questions about journalistic integrity.
@@
Figure 6. Framework for designing chatbots in online news. Our framework includes four different facets informing the policy. Once the policy is formalized, it needs to be communicated through the feedback loop. The informing and feedback loop can be iterated multiple times.
@@
Figure 6. Framework for designing chatbots in online news. Our framework includes four different facets informing the policy. Once the policy is formalized, it needs to be communicated through the feedback loop. The informing and feedback loop can be iterated multiple times.
@@ 6. A Framework for Designing Chatbots
Overall, the findings from the two studies suggest that there is a high-level agreement between journalists and readers about the role of a QA chatbot in newsrooms. Both parties predominantly wanted the chatbot to answer factual questions while they wanted to engage with each other directly for subjective questions. However, as mentioned above, our findings suggest readers do not see the role of a chatbot to be a zero-sum situation, i.e., all factual questions should be answered by the chatbot while journalists should focus on answering questions requiring subjective interpretation. A considerable amount of readers wanted the chatbot to answer factual as well as subjective questions. Similarly, we also noticed that readers want authors to answer a substantial number of factual questions, despite asking most of the factual questions to the chatbot.
Given this somewhat conflicting alignment between readers and journalists, it is clear that researchers and news organizations need a systemic approach to calibrate and communicate the needs and expectations of the two stakeholders. We propose a conceptual framework with four facets to facilitate this. For ease of explanation, we assume that a research group is developing the QA chatbot. However, other user groups (news organizations, tech industry) can use our framework as well. Finally, while our discussion below is specific to QA chatbots, we believe the framework will be useful for designing a wide range of chatbots.
@@ 6.1. Facets
We briefly describe the four main facets or components of the framework below.
@@
Technology. The final facet is understanding the design space of the currently available technology. News organizations should conduct a review such as ours in Section 2, consult internal research and developer teams, or recruit external researchers or corporate tech companies for this purpose.
@@ 6.2. Inform Stage
The two arrows from the authors and readers in Figure 6 (through the two studies) have informed us that any future QA chatbot should be able to answer factual questions. However, the arrow from technology indicates that the most powerful technology available today, LLMs, can hallucinate, provide information without proper references, or even provide wrong information (Weidinger et al., 2021). Thus, researchers should consider how they can support factual QA during this stage. For instance, instead of relying solely on the LLMs, researchers can consider a hybrid model that uses an expert user base for this purpose. Xiao et al. recently showed that coupling AI with an expert user base can support the health information needs of general users (Xiao et al., 2023). Researchers from Microsoft Research recently proposed LLM-Augmenter (Peng et al., 2023), a system that can provide answers to factual questions by using external knowledge bases. One good knowledge base for news articles could be the background research and references collected by the authors for writing the articles (Caswell, 2019; Caswell and Dörr, 2018).
@@
Our findings indicate that researchers will need to devise a routing policy for factual and subjective questions in the future. One practical solution would be to adopt the predominant view between journalists and readers: the chatbot will only answer factual questions while re-routing all questions seeking opinions and explanations to the authors. While this policy may not cater to the requirements of every reader, it will likely guarantee legal compliance for the news organization. For example, asking a chatbot to provide human-like opinion may violate legislative principles penned in the “Blueprint For An AI Bill Of Rights” (Blueprint for an AI Bill of Rights, 2023; Deshpande et al., 2023).
Researchers can adopt machine learning approaches to achieve this routing mechanism. For example, researchers can use soft-prompting (Qin and Eisner, 2021) with a pre-trained LLM to separate factual questions from subjective questions. Alternatively, researchers can collect a large corpus of questions similar to ours and then fine-tune a pre-trained LLM for this task. Finally, our findings suggest readers may expect answers to some factual questions directly from the authors, instead of the chatbot. To facilitate that, researchers can develop a scoring model that will measure the severity of a factual question and based on that route the question to either the authors or chatbot.
@@ 6.2.3. Collecting Critical Feedback from Readers
Our findings suggest overreliance on chatbots may deprive newsrooms of obtaining critical feedback from readers such as the underlying causes for the low perceived quality of an article or what raised questions about journalistic integrity in readers’ minds. According to our results, readers may not raise questions about journalistic integrity to a chatbot as frequently as they would to the authors, even if they think the quality of the article is low. Several previous works have reported similar negative impact of chatbots in user experience design (Luger and Sellen, 2016; Liao et al., 2016). Considering this fact, researchers may recommend news organizations to plan alternative methods to collect critical feedback from readers. As mentioned by journalists in the interviews, many organizations now organize audience gathering and feedback sessions, which could be an excellent alternative to achieve this.
@@ 6.3. Feedback Stage
The final part of our framework is the feedback loop. All stakeholders from the four facets should know the policy determined after the inform stage. For example, based on the discussion in Section 6.2.2., if researchers and the news organization decide that the chatbot will not provide opinions and explanations, that decision should be conveyed to the readers. To encourage readers to provide critical feedback, news organizations can inform readers that questions about journalistic integrity will be immediately rerouted to authors. News organizations should also advertise other mediums for providing feedback (e.g., member gatherings) to readers. Finally, researchers can iterate on the inform and feedback stages for finalizing the chatbot policy.
@@ 7.1.2. Increase Trust and Accountability
News organizations are grappling with a crisis of trust among their audiences. Prior research has demonstrated, presenting information alongside explanations of how and why news stories were developed can significantly enhance the trust readers place in news outlets (Masullo et al., 2017). This study adopts a novel approach by directly engaging with readers and inviting them to share their inquiries about the news. Successfully answering these questions about the news production process would quench readers thirst about news production issues (e.g., elaborating on the rationale behind including certain sources while omitting others) and would help increase trust and accountability.
@@ 7.1.3. New Revenue Stream
News industry, particularly local news is suffering from financial troubles (HOLYK, 2011; Kirchhoff, 2010). A new LLM based audience engagement could open a new revenue stream. For instance, news industry can offer the LLM based chatbot services through API to third-party applications such as Google Home, Alexa, Siri, etc. Also, readers’ questions can be utilized, with permission, for data mining and extract insights for better news recommendation and advertisement. Moreover, each newsroom has its own archive of historical news. Currently, the archives are mostly reserved for record purpose without being involved in any revenue process. LLMs can utilize this valuable news archive to answer questions and bring this resource in a revenue process.
@@ 7.2. Limitations and Future Work
Although we interviewed journalists and conducted an in-depth analysis of the readers’ questions, our analysis is not free from limitations. For example, while we engaged a significant number of readers in the study, a small-scale followup interview study with readers might reveal readers’ perspective about chatbots more clearly. We consider such a study as our immediate next step. We also did not study how the questions might differ between authors and non-authors or humans and AI. These are interesting facets that we want to study in the future. Our dataset contains several other facets that we have not explored yet. For example, we have not looked into the effects of news topics. We also have not looked into types of news such as investigative, breaking, etc. These analyses are left for future work as they were less relevant to our current research questions.
@@ 8. Conclusion
News industry is exploring how to integrate Large Language Models (LLMs) and Generative AI into the news production, distribution, audience engagement and other related processes. In this paper, we took the first step towards designing an LLM powered chatbot for addressing readers questions after reading news articles. We examined what questions do readers ask to authors and chatbot and if there are any differences. To do that, we interviewed journalists and conducted controlled experiments with readers. Our findings show that the questioning behavior of readers vary depending on who is receiving the question- author or chatbot. One of the major insights from our results is that readers ask factual questions more to chatbot compared to authors. Another significant observation is that readers question about journalistic integrity more to authors compared to the chatbot. From interviewing journalists, we observe that journalists prefer a human-in-the-loop process rather than delegating all questions to chatbot; particularly for the opinionated questions. Our findings have implications for designing a chatbot for audience engagement; specifically to evaluate how correctly LLMs can answer the questions that readers normally ask as we found. We argue that a careful and systematic approach in introducing this LLM technology to a trust-sensitive news industry would help minimize risks and ethical concerns.

### 24 Mar 2024 v2

@@ 5.3. Effect of Perceived Quality and Personal Preferences (RQ3, RQ4)
To measure how perceived quality influences questions, we adopted a structural equation model (SEM) (Ullman and Bentler, 2012), shown in Figure 5. Similar to previous research (Graefe et al., 2018), we consider perceived quality to be a latent variable and derive it from user-provided ratings on readability, credibility, and journalistic expertise. Figure 9 in Appendix A shows participants’ ratings across the three higher-level dimensions that capture the perceived quality of the articles. Overall, participants rated the articles highly across the three dimensions. To answer RQ4, we also added participants’ preferences for news outlets in the model. We manually checked the preferred news outlets provided by the participants (Figure 7e), matched the polarity of the outlets to AllSides ratings (AllSides, 2023), and then categorized them to either left, neutral, or right-leaning news outlets. We assigned neutral to participants who read both left and right-leaning news (27). We assigned a left-leaning tag if the participant consumed mostly left-leaning news (60) and right-leaning (37) vice versa. Finally, we conducted multigroup analysis (Grace and Jutila, 1999) to capture differences between the authors and chatbot condition. We used the piecewiseSEM (Lefcheck, 2016) package in R to find which paths vary across the two conditions (i.e., groups).
Table 3 presents the results from the SEM analysis. The model fit measures are: Comparative Fit Index (CFI): 0.96; Goodness of Fit (GFI): 0.90; Adjusted Goodness of Fit (AGFI): 0.91; $\chi^{2}$: 610.59; and RMSEA: 0.07. The measures indicate a good SEM model fit.. The coefficients show the polarity and strength of the relations. For example, the coefficient for Journalistic Integrity and Perceived Quality was -0.54. This means low-quality articles correlated with a high number of questions about journalistic integrity.
@@
Figure 6. Framework for designing chatbots in online news. Our framework includes five different facets informing the policy. Once the policy is formalized, it needs to be communicated through the feedback loop. The informing and feedback loop can be iterated multiple times.
@@
Figure 6. Framework for designing chatbots in online news. Our framework includes five different facets informing the policy. Once the policy is formalized, it needs to be communicated through the feedback loop. The informing and feedback loop can be iterated multiple times.
@@ 6. A Framework for Designing Chatbots
Overall, the findings from the two studies suggest that there is a high-level agreement between journalists and readers about the role of a QA chatbot in newsrooms. Both parties predominantly wanted the chatbot to answer factual questions while they wanted to engage with each other directly for subjective questions. However, as mentioned above, our findings suggest readers do not see the role of such a chatbot to be a zero-sum situation, i.e., all factual questions should be answered by the chatbot while journalists should focus on answering questions requiring subjective interpretation. A considerable amount of readers wanted the chatbot to answer factual as well as subjective questions. Similarly, we also noticed that readers want authors to answer a substantial number of factual questions, despite asking most of the factual questions to the chatbot.
Given this somewhat conflicting alignment between readers and journalists, it is clear that researchers and news organizations need a systemic approach to calibrate and communicate the needs and expectations of the two stakeholders. We propose a conceptual framework with four facets to facilitate this. For ease of explanation, we assume that a newsroom or news organization is developing the QA chatbot. However, other user groups (researchers, tech industry) can use our framework as well. Finally, while our discussion below is specific to QA chatbots, we believe the framework will be useful for designing a wide range of chatbots.
@@ 6.1. Facets
We briefly describe the five main facets or components of the framework below.
@@
Technology. The fourth facet is understanding the design space of the currently available technology. News organizations should conduct a review such as ours in Section 2, consult internal research and developer teams, or recruit external researchers or corporate tech companies for this purpose.
AI Regulatory Policies. The final facet is seeking legal consults to understand current AI/technology regulatory policies such as the “General Data Protection Regulation (GDPR)” (Voigt and Von dem Bussche, 2017) from the European Union (EU) or the “Blueprint For An AI Bill Of Rights” (Blueprint for an AI Bill of Rights, 2023) from the White House in the U.S. These policies may contain rules and policies that news organizations should abide by while designing the chatbot.
@@ 6.2. Inform Stage
The two arrows from the authors and readers in Figure 6 (through the two studies) have informed us that any future QA chatbot should be able to answer factual questions. However, the arrow from technology indicates that the most powerful technology available today, LLMs, can hallucinate, provide information without proper references, or even provide wrong information (Weidinger et al., 2021). Thus, newsrooms should consider how they can support factual QA during this stage. For instance, instead of relying solely on the LLMs, researchers can consider a hybrid model that uses an expert user base for this purpose. Xiao et al. recently showed that coupling AI with an expert user base can support the health information needs of general users (Xiao et al., 2023). Researchers from Microsoft Research recently proposed LLM-Augmenter (Peng et al., 2023), a system that can provide answers to factual questions by using external knowledge bases. One good knowledge base for news articles could be the background research and references collected by the authors for writing the articles (Caswell, 2019; Caswell and Dörr, 2018).
@@
Our findings indicate that researchers will need to devise a routing policy for factual and subjective questions in the future. One practical solution would be to adopt the predominant view between journalists and readers: the chatbot will only answer factual questions while re-routing all questions seeking opinions and explanations to the authors. While this policy may not cater to the requirements of every reader, it will likely guarantee legal compliance for the news organization. For example, the AI regulatory policy facet in the framework can inform the newsroom that asking a chatbot to provide human-like opinion may violate legislative principles penned in the “Blueprint For An AI Bill Of Rights” (Blueprint for an AI Bill of Rights, 2023; Deshpande et al., 2023).
Newsrooms can adopt machine learning approaches to achieve this routing mechanism. For example, newsrooms can use soft-prompting (Qin and Eisner, 2021) with a pre-trained LLM to separate factual questions from subjective questions. Alternatively, newsrooms can collect a large corpus of questions similar to ours and then fine-tune a pre-trained LLM for this task. Finally, our findings suggest readers may expect answers to some factual questions directly from the authors, instead of the chatbot. To facilitate that, newsrooms can develop a scoring model that will measure the severity of a factual question and based on that route the question to either the authors or chatbot.
@@ 6.2.3. Collecting Critical Feedback from Readers
Our findings suggest that over-reliance on chatbots may deprive newsrooms of obtaining critical feedback from readers such as the underlying causes for the low perceived quality of an article or what raised questions about journalistic integrity in readers’ minds. According to our results, readers may not raise questions about journalistic integrity to a chatbot as frequently as they would to the authors, even if they think the quality of the article is low. Several previous works have reported similar negative impacts of chatbots in user experience design (Luger and Sellen, 2016; Liao et al., 2016). Considering this fact, researchers may recommend news organizations to plan alternative methods to collect critical feedback from readers. As mentioned by journalists in the interviews, many organizations now organize audience gathering and feedback sessions, which could be an excellent alternative to achieve this.
@@ 6.3. Feedback Stage
The final part of our framework is the feedback loop. All stakeholders from the five facets should know the policy determined after the inform stage. For example, based on the discussion in Section 6.2.2., if the news organization decides that the chatbot will not provide opinions and explanations, that decision should be conveyed to the readers. To encourage readers to provide critical feedback, news organizations can inform readers that questions about journalistic integrity will be immediately rerouted to authors. News organizations should also advertise other mediums for providing feedback (e.g., member gatherings) to readers. Finally, the news organization can iterate on the two stages for finalizing the chatbot policy.
@@ 7.1.2. Increase Trust and Accountability
News organizations are grappling with a crisis of trust among their audiences. Prior research has demonstrated that presenting information alongside explanations of how and why news stories were developed can significantly enhance the trust readers place in news outlets (Masullo et al., 2017). This study adopts a novel approach by directly engaging with readers and inviting them to share their inquiries about the news. Successfully answering these questions about the news production process would quench readers’ thirst for news production issues (e.g., elaborating on the rationale behind including certain sources while omitting others) and would help increase trust and accountability.
@@ 7.1.3. New Revenue Stream
The news industry, particularly local news is suffering from financial troubles (HOLYK, 2011; Kirchhoff, 2010). A new LLM-based audience engagement could open a new revenue stream. For instance, the news industry can offer LLM-based chatbot services through API to third-party applications such as Google Home, Alexa, Siri, etc. Also, readers’ questions can be utilized, with permission, for data mining and extracting insights for better news recommendations and advertisements. Moreover, each newsroom has its own archive of historical news. Currently, the archives are mostly reserved for record purposes without being involved in any revenue process. LLMs can utilize this valuable news archive to answer questions and bring this resource into a revenue process.
@@ 7.2. Limitations and Future Work
Although we interviewed journalists and conducted an in-depth analysis of the readers’ questions, our analysis is not free from limitations. For example, while we engaged a significant number of readers in the study, a small-scale followup interview study with readers might reveal readers’ perspective about chatbots more clearly. We consider such a study as our immediate next step. Our dataset contains several other facets that we have not explored yet. For example, we have not looked into the effects of news topics. We also have not looked into types of news such as investigative, breaking, etc. These analyses are left for future work as they were less relevant to our current research questions.
@@
The study might also involve hidden confounding variables. For instance, there are two differences between the author and chatbot conditions: human vs AI and authors vs non-authors. We did not study these facets separately as our goal was to understand how readers direct questions towards authors and chatbot, not how the questions differ between human vs AI or authors vs non-authors. In other words, we were interested in the complete constructs embodied by authors and chatbot, not their individual components. Nevertheless, we acknowledge that studying these facets could provide more insights into readers’ questioning patterns.
@@ 8. Conclusion
The news industry is exploring how to integrate Large Language Models (LLMs) and Generative AI into news production, distribution, audience engagement, and other related processes. In this paper, we took the first step towards designing an LLM-powered chatbot by understanding how journalists and readers want to use such a chatbot. To achieve that, we interviewed journalists and conducted controlled experiments with readers. From interviewing journalists, we observe that journalists prefer a human-in-the-loop process rather than delegating all questions to a chatbot; particularly for the opinionated questions. We then examined what questions readers ask the authors and chatbot and if there are any differences. Our findings show that the questioning behavior of readers varies depending on who is receiving the question- the author or the chatbot. One of the major insights from our results is that readers ask factual questions more to chatbots compared to authors. Another significant observation is that readers question about journalistic integrity more to authors compared to the chatbot. Our findings have implications for designing a chatbot for audience engagement; specifically to evaluate how correctly LLMs can answer the questions that readers normally ask as we found. We argue that a careful and systematic approach to introducing this LLM technology to a trust-sensitive news industry would help minimize risks and ethical concerns.